2Sin a=2-Cos A, FIND SinA
 

Grade:10

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hi



please check yourself too.

Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty
Rinkoo Gupta
9 years ago

Cosa=2-2sina=2(1-sina)

Squaring, we get

Cos2a=4(1-sina)2

1-sin2a=4(1+sin2a-2sina)

5sin2a-8sina+3=0

Solving we get sina=1,3/5

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty